OpenStudy (anonymous):

Factor out please: 4x^2 + 24x+ 36

mathslover (mathslover):

let me explain this to you : \(\huge{a^2+2(a)(b)+b^2=(a+b)^2}\)---------FORMULA \[\huge{(2x)^2+2(2x)(6)+6^2}\] compare this equation with the formula a =2x and b = 6 hence \(\huge{4x^2+24x+36=(2x+6)^2}\)